OpenStudy (anonymous):

Describe how to find an equation of the line through the point (-2, 1) with slope 3/2.

OpenStudy (anonymous):

SLOPE: Y(2) - Y(1)/X(2) - X(1) ----> THE NUMBERS INSIDE THE "( )" ARE SUBSCRIPT Y(2) - Y(1)/X(2) - X(1) = 3/2 Y(2) - 1/X(2) - (-2) = 3/2 Y = 3X/2 + 4

OpenStudy (anonymous):

did you get it???? :-)

OpenStudy (anonymous):

The equation is y=3x/6??? I'm confused. LOL

OpenStudy (anonymous):